# Tensor Attention Training: Provably Efficient Learning of Higher-order Transformers

### v1

@@ 3.2 Key Definitions of Tensor Attention
Let $\epsilon,B>0$. Let $A_{1},A_{2},A_{3},A_{4},A_{5},E\in\mathbb{R}^{n\times d}$ and let $X_{1},X_{2},X_{3},Y_{1},Y_{2}\in\mathbb{R}^{d\times d}$ (see Definition 3.10). Let $X=X_{1}\cdot(X_{2}\oslash X_{3})^{\top}\in\mathbb{R}^{d\times d^{2}}$. Assume that $\max\{\|A_{1}X_{1}\|_{\infty}$, $\|A_{2}X_{2}\|_{\infty}$, $\|A_{3}X_{3}\|_{\infty}$, $\|A_{4}Y_{1}\|_{\infty}$, $\|A_{5}Y_{2}\|_{\infty}$, }$\}\leq B$. Let us assume that any numbers in the previous matrices are in the $\log(n)$ bits model. We define $\mathsf{Loss}(X)$ the same as Definition 3.10. Let the gradient of loss function $\mathsf{Loss}(X)$ be $\frac{\mathrm{d}\mathsf{Loss}(X)}{\mathrm{d}X}\in\mathbb{R}^{d\times d^{2}}$. Then, our target is to output a matrix $\widetilde{g}\in\mathbb{R}^{d\times d^{2}}$ satisfying
$\|\widetilde{g}-\frac{\mathrm{d}\mathsf{Loss}(X)}{\mathrm{d}X}\|_{\infty}\leq\epsilon.$
@@ 4 Exact Tensor Attention Gradient Computation and Complexity
In this section, we provide the closed form of the tensor attention gradient of the loss function (Definition 3.10) and also its computational time. First, we can calculate the closed form of the gradient in the following lemma, whose proof is in Appendix E.5.
Lemma 4.1 (Close form of gradient, informal version of Lemma E.6).
Define the function $\mathsf{F}(x)\in\mathbb{R}^{n\times n^{2}}$ as in Definition D.6. Suppose that $A_{1},A_{2},A_{3}\in\mathbb{R}^{n\times d}$ are three given matrices. Suppose that $\mathsf{Loss}(x)$ is defined as Definition 3.10, where $x=\operatorname{vec}(X)$. Then, we can show
@@
Suppose that $A_{1},A_{2},A_{3},$ $A_{4},A_{5},E\in\mathbb{R}^{n\times d}$ are input fixed matrices. We denote matrix variables as $X\in\mathbb{R}^{d\times d^{2}}$ and $Y\in\mathbb{R}^{d^{2}\times d}$ (gradient computation is w.r.t. $X$ ). Let $g=\frac{\mathrm{d}\mathsf{Loss}(X)}{\mathrm{d}X}\in\mathbb{R}^{d\times d^{2}}$ (for definition of $\mathsf{Loss}(X)$, see Definition 3.10). Then we can demonstrate that computing the gradient $g\in\mathbb{R}^{d\times d^{2}}$ requires ${\cal T}_{\mathrm{mat}}(n,d^{2},n^{2})$ time.
Note that ${\cal T}_{\mathrm{mat}}(n,d^{2},n^{2})\geq\Omega(n^{3})$. Thus, the naive tensor attention gradient computation is a complexity obstacle in practice as discussed in Section 1. Based on the closed formulation in Lemma 4.1, we will have our acceleration method, which will be introduced in the following section.
5 Fast Gradient Computation via Tensor Trick and Polynomial Method
In this section, we show how to compute the tensor attention matrix gradient in almost linear time. In Section 5.1, we demonstrate our main results. In Section 5.2, we introduce some key tensor tricks that are used in our proof.
@@ 5.1 Main Results for Fast Gradient Computation
Based on the polynomial approximation methods, [21] get the following lemma about tensor attention acceleration, which will be used in our main results proof.
@@
Using similar polynomial approximation methods, with many more tensor analysis techniques (Section 5.2), we can get our main acceleration results.
@@
We can use the polynomial approximation method to get low-rank approximation results for $D^{-1}\exp(A_{1}X(A_{2}\otimes A_{3})^{\top}/d)$ in Lemma F.1. Then, based on the close form of the tensor attention gradient solution in Theorem 4.3 and many tensor tricks (Section 5.2 and Appendix C), we can smartly convey these low rank properties to the gradient formulation, where two key steps need to be fixed (see details in Lemma F.5 and Lemma F.7). ∎
@@
The assumption in Theorem 5.2 is practical. In practice, the $n$ is super large, e.g, $n=2,000,000$ for Google’s Gemini 1.5 Pro  [48], while the model training uses a half-precision floating-point format, e.g., the bit number is a constant $16$. On the other hand, our assumption is “tight”, where if we slightly weaken the assumption, there is no algorithm that can solve the tensor attention gradient computation in truly sub-cubic complexity (Theorem 6.3).
Our Theorem 5.2 can accurately approximate ($\epsilon=1/\operatorname{poly}(n)$) the tensor attention gradient computation in almost linear time $n^{1+o(1)}$ under practical assumptions (see discussion in Remark 5.3). Thus, our methods solve the last puzzle of tensor attention acceleration. This may make tensor attention practical, as we can get around the cubic time complexity barrier both in inference and training.
@@ 5.2 Tensor Operation Analysis Techniques
Here, we introduce some key techniques during proof Theorem 5.2. These techniques make it possible to convey the low-rank property even during the tensor computation.
We first introduce a swap rule and a distributed rule, where both proofs are in Appendix C.2
Fact 5.4 (Swap rule for tensor product and matrix product).
@@
Fact 5.4 tells us that we can swap the order of tensor operation and matrix multiplication, so we can always compute the low dimension first to reduce the complexity.
@@
Fact 5.5 tells us that the multiple tensor operation can be distributed to a different format. If we have some low-rank matrix/tensor, we can distribute them to each component so that each component can be accelerated. The high-level intuition sounds like we borrow some benefits from other terms to fix the bottleneck term.
We prove an important tool whose proof is in Appendix C.2.
@@
Given $A_{1}\in\mathbb{R}^{n_{1}\times d_{1}}$, $A_{2}\in\mathbb{R}^{n_{2}\times d_{1}}$, let $A:=(A_{1}\oslash A_{2})\in\mathbb{R}^{n_{1}n_{2}\times d_{1}}$. Given $B_{1}\in\mathbb{R}^{n_{1}\times d_{2}}$, $B_{2}\in\mathbb{R}^{n_{2}\times d_{2}}$, let $B:=(B_{1}\oslash B_{2})\in\mathbb{R}^{n_{1}n_{2}\times d_{2}}$. We define $C\in\mathbb{R}^{d_{1}\times d_{2}}$ as $C:=A^{\top}B$ and $C_{1}:=A_{1}^{\top}B_{1}\in\mathbb{R}^{d_{1}\times d_{2}},C_{2}:=A_{2}^{\top}B_{2}\in\mathbb{R}^{d_{1}\times d_{2}}$. Then, we have
  • Part 1. $C_{1}\circ C_{2}=C$
  • Part 2. Given $A_{1},A_{2},B_{1},B_{2}$, we can get $C$ in ${\cal T}_{\mathrm{mat}}(d_{1},\max\{n_{1},n_{2}\},d_{2})$ time.
Lemma C.14 is a highly non-trivial method to handle tensor operation, $\circ$ and matrix multiplication together. By using the method, we can save the computation time from ${\cal T}_{\mathrm{mat}}(d,n^{2},d)$ to ${\cal T}_{\mathrm{mat}}(d,n,d)$, which gets rid of the bottleneck quadratic term $n^{2}$.
@@
6 Tensor Attention Gradient Computation Complexity Lower Bound
In this section, we will show that our assumption is “tight”. First, we introduce some hardness analysis background in Section 6.1. Then, we will introduce our main hardness result in Section 6.2.
6.1 Strong Exponential Time Hypothesis and Tensor Attention Forward Hardness
We provide the findings that our results are based on. We first introduce a well-known hypothesis. The Strong Exponential Time Hypothesis ($\mathsf{SETH}$), a well-established conjecture, has been instrumental in establishing fine-grained lower bounds for numerous algorithmic problems, as highlighted in the survey by [133]. More than two decades ago, [61] introduced $\mathsf{SETH}$ as an enhanced version of the $\mathsf{P}\neq\mathsf{NP}$ conjecture, positing that current solving $\mathsf{SAT}$ problem algorithms are nearly optimal in terms of efficiency.
@@ 6.1 Strong Exponential Time Hypothesis and Tensor Attention Forward Hardness
We will critically utilize the hardness result of the forward tensor attention computation from previous work as follows.
@@
Assuming $\mathsf{SETH}$, for any constant $\delta>0$, no algorithm can solve $\mathsf{ATAttC}(n,d=\Theta(\log n),B=\Theta(\sqrt[3]{(1+\gamma)\log n}),\epsilon=n^{\gamma-O(1)})$ (Definition 3.9) in $O(n^{3-\delta})$ time, even if the inputs meet the following conditions for any $\gamma\geq 0$: (1) $V\in\{0,1\}^{n^{2}\times d}$, (2) There exists $B_{a}\leq O((1+\gamma)\log^{2}n)=O(d(\sqrt[3]{(1+\gamma)\log n})^{3})$ where all entries of $QK^{\top}$ are within the range $[1,B_{a}]$ and more than half entries in each row of $QK^{\top}$ are equal to $B_{a}$.
If we change the assumption from $B=O(\sqrt[3]{\log n})$ to $B=\Theta(\sqrt[3]{(1+\gamma)\log n}$ with $\gamma=\omega(1)$, Lemma 6.2 shows that tenor attention forward computation is hard, i.e., no algorithm can solve it in truly sub-cubic time (assume $\mathsf{SETH}$).
6.2 Main Results for Hardness
Based on the above observation (Lemma 6.2), we can prove our main results for tensor attention gradient computation hardness.

### v2

@@ 3.2 Key Definitions of Tensor Attention
Let $\epsilon,B>0$. Let $A_{1},A_{2},A_{3},A_{4},A_{5},E\in\mathbb{R}^{n\times d}$ and let $X_{1},X_{2},X_{3},Y_{1},Y_{2}\in\mathbb{R}^{d\times d}$ (see Definition 3.8). Let $X=X_{1}\cdot(X_{2}\oslash X_{3})^{\top}\in\mathbb{R}^{d\times d^{2}}$. Assume that $\max\{\|A_{1}X_{1}\|_{\infty}$, $\|A_{2}X_{2}\|_{\infty}$, $\|A_{3}X_{3}\|_{\infty}$, $\|A_{4}Y_{1}\|_{\infty}$, $\|A_{5}Y_{2}\|_{\infty}\}\leq B$. Let us assume that any numbers in the previous matrices are in the $\log(n)$ bits model. We define $\mathsf{Loss}(X)$ the same as Definition 3.8. Let the gradient of loss function $\mathsf{Loss}(X)$ be $\frac{\mathrm{d}\mathsf{Loss}(X)}{\mathrm{d}X}\in\mathbb{R}^{d\times d^{2}}$. Then, our target is to output a matrix $\widetilde{g}\in\mathbb{R}^{d\times d^{2}}$ satisfying $\|\widetilde{g}-\frac{\mathrm{d}\mathsf{Loss}(X)}{\mathrm{d}X}\|_{\infty}\leq\epsilon.$
@@ 4 Exact Tensor Attention Gradient Computation and Complexity
In this section, we provide the closed form of the tensor attention gradient of the loss function (Definition 3.8) and also its computational time. First, we calculate the closed form of the gradient in the following lemma, whose proof is in Appendix E.5.
Lemma 4.1 (Closed form of gradient, informal version of Lemma E.6).
Define the function $\mathsf{F}(x)\in\mathbb{R}^{n\times n^{2}}$ as in Definition D.6. Suppose that $A_{1},A_{2},A_{3}\in\mathbb{R}^{n\times d}$ are three given matrices. Suppose that $\mathsf{Loss}(x)$ is defined as Definition 3.8, where $x=\operatorname{vec}(X)$. Then, we have
@@
Suppose that $A_{1},A_{2},A_{3},$ $A_{4},A_{5},E\in\mathbb{R}^{n\times d}$ are input fixed matrices. We denote matrix variables as $X\in\mathbb{R}^{d\times d^{2}}$ and $Y\in\mathbb{R}^{d^{2}\times d}$ (gradient computation is w.r.t. $X$ ). Let $g=\frac{\mathrm{d}\mathsf{Loss}(X)}{\mathrm{d}X}\in\mathbb{R}^{d\times d^{2}}$ (for definition of $\mathsf{Loss}(X)$, see Definition 3.8). Then, we show that computing the gradient $g\in\mathbb{R}^{d\times d^{2}}$ requires ${\cal T}_{\mathrm{mat}}(n,d^{2},n^{2})$ time.
Note that ${\cal T}_{\mathrm{mat}}(n,d^{2},n^{2})\geq\Omega(n^{3})$. Thus, the naive tensor attention gradient computation is a complexity obstacle in practice as discussed in Section 1. Based on the closed formulation in Lemma 4.1, we derive our acceleration method, which will be introduced in the following section.
5 Fast Tensor Attention Gradient Computation
In this section, we show how to compute the tensor attention matrix gradient in almost linear time. In Section 5.1, we demonstrate our main results. In Section 5.2, we introduce some key tensor techniques used in our proof.
Algorithm 1 Almost Linear Time Tensor Attention Gradient Computation
1:procedure FastTensorAttention($A_{1},A_{2},A_{3},A_{4},A_{5},E\in\mathbb{R}^{n\times d},X_{1},X_{2},X_{3},Y_{1},Y_{2}\in\mathbb{R}^{d\times d},n\in\mathbb{N}_{+},d\in\mathbb{N}_{+},\epsilon\in(0,0.1)$) $\triangleright$ Definition 3.9, Theorem 5.2
2:     $\triangleright$ $n$ can be viewed as the length of the sentence
3:     $\triangleright$ $d$ can be viewed as the feature of dimension, and we assume $d=O(\log n)$
4:     $\triangleright$ $\epsilon$ is the accuracy output, and we typically pick $1/\operatorname{poly}(n)$
5:     Get $U_{1},V_{1},W_{1}\in\mathbb{R}^{n\times n^{o(1)}}$     to approximate $\mathsf{S}(x)$ via Lemma F.1 $\triangleright$ $O(n^{1+o(1)})$ time
6:     $U_{2}\leftarrow U_{1}(V_{1}\oslash W_{1})^{\top}\mathsf{L}(y)-E$     to approximate $\mathsf{V}(x)$ via Lemma 5.6 $\triangleright$ $O(n^{1+o(1)})$ time
7:     $V_{2},~{}W_{2}\leftarrow A_{4}Y_{1},~{}A_{5}Y_{2}$     to approximate $\mathsf{W}(x)$ via Lemma F.3 $\triangleright$ $O(nd^{2})$ time
8:     $U_{3},~{}V_{3},~{}W_{3}\leftarrow U_{1}\ominus U_{2},~{}V_{1}\ominus V_{2},~{}W_{1}\ominus W_{2}$     to approximate $\mathsf{F}_{a}(x)$ via Lemma F.5 $\triangleright$ $O(n^{1+o(1)})$ time
9:     Precompute $V_{1}^{\top}V_{2}$ and $W_{1}^{\top}W_{2}$     to approximate $\mathsf{F}_{b}(x)$ via Lemma F.7 $\triangleright$ $O(n^{1+o(1)})$ time
10:     for $j_{0}\in[n]$ do $\triangleright$ Overall $\widetilde{\mathsf{R}}(x)$ takes $O(n^{1+o(1)})$ time
11:         $\widetilde{\mathsf{R}}(x)_{j_{0}}\leftarrow(U_{1})_{j_{0},*}((V_{1}^{\top}V_{2})\circ(W_{1}^{\top}W_{2}))((U_{2})_{j_{0},*})^{\top}$
12:     end for
13:     $U_{4}\leftarrow\operatorname{diag}(\widetilde{\mathsf{R}}(x))U_{1}$triangleright $O(n^{1+o(1)})$ time
14:     $V_{4},~{}W_{4}\leftarrow V_{1},~{}W_{1}$triangleright $O(n^{1+o(1)})$ time
15:      /* Approximate $\mathsf{F}(x)$, Theorem F.8 */
16:     $U_{5},~{}V_{5},~{}W_{5}\leftarrow[U_{3},-U_{4}],~{}[V_{3},V_{4}],~{}[W_{3},W_{4}]$triangleright $O(n^{1+o(1)})$ time
17:      /* Approximate $g$, Theorem F.8 */
18:     Precompute $A_{1}^{\top}U_{5}$, $A_{2}^{\top}V_{5}$, $A_{3}^{\top}W_{5}$ separately $\triangleright$ $O(dn^{1+o(1)})$ time
19:     $\widetilde{g}\leftarrow(A_{1}^{\top}U_{5})\odot(A_{2}^{\top}V_{5})\odot(A_{3}^{\top}W_{5})$triangleright $\odot$ in Definition C.3. $O(d^{3}n^{o(1)})$ time
20:     return $\widetilde{g}$triangleright As $d=O(\log n)$, the total complexity is $O(n^{1+o(1)})$ time
21:end procedure
@@ 5.1 Main Results for Fast Gradient Computation
Based on the polynomial approximation methods, [15] get the following result about tensor attention acceleration, which will be used to prove our main result.
@@
Using similar polynomial approximation methods, and combined with a series of tensor analysis techniques (Section 5.2), we get our main acceleration results.
@@
Assuming the entries of $A_{1},A_{2},$ $A_{3},$ $A_{4},A_{5},E\in\mathbb{R}^{n\times d}$ and $X_{1},X_{2},X_{3},Y_{1},Y_{2}\in\mathbb{R}^{d\times d}$ are represented using $O(\log n)$ bits. Then, there exist an algorithm (Algorithm 1) that runs in $n^{1+o(1)}$ time to solve $\mathsf{ATAttLGC}(n,d=O(\log n),B=o(\sqrt[3]{\log n}),\epsilon=1/\operatorname{poly}(n))$ (see Definition 3.9), i.e., our algorithm computes a gradient matrix $\widetilde{g}\in\mathbb{R}^{d\times d^{2}}$ satisfying $\|\frac{\mathrm{d}\mathsf{Loss}(X)}{\mathrm{d}X}-\widetilde{g}\|_{\infty}\leq 1/\operatorname{poly}(n)$.
@@
We use the polynomial approximation method to obtain low-rank approximation results for $D^{-1}\exp(A_{1}X(A_{2}\otimes A_{3})^{\top}/d)$ in Lemma F.1. However, this cannot be directly used for the closed form of the tensor attention gradient solution in Theorem 4.3. Utilizing a series of tensor techniques (Section 5.2 and Appendix C), we smartly convey these low rank properties throughout the gradient formulation and computation, where two key steps are fixed in Lemma F.5 and Lemma F.7. ∎
@@
The assumption in Theorem 5.2 is practical. In practice, especially in recent long context tasks, the $n$ is large, e.g, $n=2\times 10^{6}$ for Google’s Gemini 1.5 Pro  [35], while the model training uses a half-precision floating-point format, e.g., the bit number is $16$. Furthermore, our assumption is “tight”, where if we slightly weaken the assumption, there is no algorithm that can solve the tensor attention gradient computation in truly sub-cubic complexity (Theorem 6.3).
Our Theorem 5.2 accurately approximates ($\epsilon=1/\operatorname{poly}(n)$) the tensor attention gradient computation in almost linear time $n^{1+o(1)}$ under practical assumptions (see the above Remark 5.3). Thus, our methods solve the last puzzle of tensor attention acceleration. Combined with previous work on tensor attention inference, this may make tensor attention practical, as we overcome the theoretical cubic time complexity barrier both in inference and training.
We provide Algorithm 1 for our almost linear time tensor attention training method. In the detailed algorithm, first, we construct $U_{1},V_{1},W_{1}$ in Lemma F.1. Then, we construct $U_{2},V_{2},W_{2}$ in Lemma F.3 and $U_{3},V_{3},W_{3}$ in Lemma F.5. We show how to construct $U_{4},V_{4},W_{4}$ in Lemma F.7. Finally, we construct $U_{5},V_{5},W_{5}$ and compute the gradient $g$ in almost linear time in Theorem F.8.
@@ 5.2 Tensor Operation Analysis Techniques
Here, we introduce some key techniques for proving Theorem 5.2. These techniques make it possible to convey the low-rank property even during the tensor operations, solving the novel technical challenges in tensor attention gradient computation.
We first introduce a swap rule and a distributed rule, where both proofs are in Appendix C.2.
Fact 5.4 (Swap rule for tensor and matrix product).
Let $W_{1},W_{2}\in\mathbb{R}^{d\times d},A_{1},A_{2}\in\mathbb{R}^{n\times d}$. We have
@@
Fact 5.4 tells us that we can swap the order of tensor operation and matrix multiplication, allowing us to always compute the low dimension first to reduce the complexity.
@@
Fact 5.5 tells us that the multiple tensor operation can be distributed to a different format. If we have some low-rank matrix/tensor, we can distribute them into each component, so that each component can be accelerated via the low-rank property. Intuitively, this allows us to borrow some low-rank benefits from other terms to fix the bottleneck terms.
We provide an important tool whose proof is in Appendix C.2.
@@
Given $A_{1}\in\mathbb{R}^{n_{1}\times d_{1}}$, $A_{2}\in\mathbb{R}^{n_{2}\times d_{1}}$, let $A:=(A_{1}\oslash A_{2})\in\mathbb{R}^{n_{1}n_{2}\times d_{1}}$. Given $B_{1}\in\mathbb{R}^{n_{1}\times d_{2}}$, $B_{2}\in\mathbb{R}^{n_{2}\times d_{2}}$, let $B:=(B_{1}\oslash B_{2})\in\mathbb{R}^{n_{1}n_{2}\times d_{2}}$. We define $C\in\mathbb{R}^{d_{1}\times d_{2}}$ as $C:=A^{\top}B$ and $C_{1}:=A_{1}^{\top}B_{1}\in\mathbb{R}^{d_{1}\times d_{2}},C_{2}:=A_{2}^{\top}B_{2}\in\mathbb{R}^{d_{1}\times d_{2}}$. Then, we have $C_{1}\circ C_{2}=C$ and given $A_{1},A_{2},B_{1},B_{2}$, we can get $C$ in ${\cal T}_{\mathrm{mat}}(d_{1},\max\{n_{1},n_{2}\},d_{2})$ time.
Lemma 5.6 is a highly non-trivial method to handle tensor operation, $\circ$ and matrix multiplication together. By using the method, we save the computation time from ${\cal T}_{\mathrm{mat}}(d,n^{2},d)$ to ${\cal T}_{\mathrm{mat}}(d,n,d)$, which gets rid of the bottleneck quadratic term $n^{2}$.
@@
Technical novelty over previous works. We generalize beyond the results of AS24b [15], which only provide methods for tensor attention forward. Our paper presents a detailed analysis for tensor attention backward, providing both upper bound and lower bound. Though we build on some results from AS24b [15] and AS24a [14], generalizing to tensor attention backward posed many technical challenges. These challenges are unique to our setting and not presented in previous settings like matrix attention [13, 14] or tensor attention forward [15]. To be more specific, we prove many key properties for tensor operation needed for backward though not needed for forward, including Facts 5.4 (swap rule for tensor and matrix product), 5.5 (distribution rule for tensor and matrix product), C.11 (tensor computation reduction to matrix product), C.12 (distribution rule for tensor computation), Claim C.20 (tensor product to matrix product). Fact 5.4, used as a key part to prove Lemmas F.1 and F.3, gives the swap rule for tensor operations. Lemma 5.6 supports the proof of Fact 5.5 and helps bypass the $O(n^{3}d^{2})$ time complexity bottleneck in the fast computation of $U_{2}$. Fact 5.5, crucial in proving Lemma F.5, shows the distributive nature of tensor operations. Using Facts C.11, C.12, and Claim C.20, we leverage the structure of low-rank matrices $U_{5},V_{5},W_{5}$ to prove Theorem 5.2.
6 Tensor Attention Gradient Complexity Lower Bound
In this section, we show that our assumption is necessary. First, we introduce some hardness analysis background in Section 6.1. Then, we introduce our main hardness result in Section 6.2.
6.1 $\mathsf{SETH}$ and Tensor Attention Forward Hardness
We provide the findings that our results are based on. We first introduce a well-known hypothesis in hardness analysis. The Strong Exponential Time Hypothesis ($\mathsf{SETH}$), a well-established conjecture, has been instrumental in establishing fine-grained lower bounds for numerous algorithmic problems, as highlighted in the survey by [114]. More than two decades ago, [46] introduced $\mathsf{SETH}$ as an enhanced version of the well-known $\mathsf{P}\neq\mathsf{NP}$ conjecture, positing that current algorithms solving the $\mathsf{SAT}$ problem are nearly optimal in terms of efficiency.
@@ 6.1 $\mathsf{SETH}$ and Tensor Attention Forward Hardness
We will critically utilize the hardness result of the forward tensor attention computation.
@@
Assuming $\mathsf{SETH}$, for any constant $\delta>0$, no algorithm can solve $\mathsf{ATAttC}(n,d=\Theta(\log n),B=\Theta(\sqrt[3]{(1+\gamma)\log n}),\epsilon=n^{\gamma-O(1)})$ (Definition 3.7) in $O(n^{3-\delta})$ time, even if the inputs meet the following conditions for any $\gamma\geq 0$: (1) $V\in\{0,1\}^{n^{2}\times d}$, (2) There exists $B_{a}\leq O((1+\gamma)\log^{2}n)=O(d(\sqrt[3]{(1+\gamma)\log n})^{3})$ where all entries of $Q(K_{1}\oslash K_{2})^{\top}$ are within the range $[1,B_{a}]$ and more than half entries in each row of $Q(K_{1}\oslash K_{2})^{\top}$ are equal to $B_{a}$.
This result shows that assuming $\mathsf{SETH}$, if we just slightly weaken the assumption from $B=O(\sqrt[3]{\log n})$ to $B=\Theta(\sqrt[3]{(1+\gamma)\log n})$ with $\gamma=\omega(1)$, then the tenor attention forward computation is hard, i.e., no algorithm can solve it in truly sub-cubic time.
6.2 Main Result for Hardness
Based on the above observation (Lemma 6.2), we prove our main result for tensor attention gradient computation hardness.